# Bounds of Validity for Bifurcations of Equilibria in a Class of Networked Dynamical Systems

### v2

@@ II-C Bounds of Validity for Lyapunov–Schmidt reduction
A key limitation of Lyapunov–Schmidt reduction is that the equivalence between the reduced bifurcation equations and the equilibria of the original high-dimensional system holds only locally. Consequently, the analysis of bifurcations leveraging this reduction method is valid only in a small neighborhood of the singular point. Understanding the extent of these neighborhoods is crucial for evaluating the robustness of conclusions drawn from analysis of the reduced equations. To address this limitation, in previous work we derived a general bound of validity for Lyapunov–Schmidt reduction in systems of ordinary differential equations [gupta2024estimates]. In this section we summarize our recent result.
In the context of Lyapunov–Schmidt reduction, we obtain computable lower bound approximations for the neighborhoods $\mathcal{N}\big((\alpha_{0},p^{\star})\big)$, $\mathcal{N}(\beta_{0})$ as open balls $\mathcal{B}\big((\alpha_{0},p^{\star}),r_{\parallel}\big)$ and $\mathcal{B}(\beta_{0},r_{\perp})$. The radii $r_{\parallel}$ and $r_{\perp}$ bound the regions around the singular point $(\alpha,\beta,p)=(\alpha_{0},\beta_{0},p^{\star})$ where the reduced and full bifurcation diagrams coincide topologically. In particular $r_{\parallel}$ defines the extent of equivalence along the critical subspace, i.e. the product between $\ker(J)$ and the parameter space $\Lambda$. Analogously, $r_{\perp}$ defines the extent of equivalence along $\ker(J)^{\perp}$. The algorithm to explicitly determine the bounding radii $r_{\parallel}$ and $r_{\perp}$ is described below.
@@
$M_{\perp}$ $:=\left\lVert{\left(W^{\top}D_{x}\Phi(x^{\star},p^{\star})W\right)^{-1}}\right\rVert.$ (6b)
@@
$\xi_{\perp}(\alpha,\beta,p)$ $:=W^{\top}\Bigl(D_{x}\Phi\bigl(\Gamma(\alpha,\beta),p\bigr)-D_{x}\Phi(x^{\star},p^{\star})\Bigr)W,$ (7b)
where $\xi_{\parallel p}(\alpha,p)=D_{p}\Phi\bigl(\Gamma(\alpha,\beta_{0}),p\bigr)-D_{p}\Phi(x^{\star},p^{\star})$ and $\xi_{\parallel\alpha}(\alpha,p)=D_{x}\Phi\bigl(\Gamma(\alpha,\beta_{0}),p\bigr)V$. The maps (7) quantify first-order variations of the Jacobian $D_{x}\Phi$ and parameter derivative $D_{p}\Phi$ along the critical subspace and its orthogonal complement. We use $\mathcal{B}_{\parallel}(r_{\parallel}):=\mathcal{B}\big((\alpha_{0},p^{\star}),r_{\parallel}\big)\subset\mathbb{R}^{q}\times\Lambda$ and $\mathcal{B}_{\perp}(r_{\perp}):=\mathcal{B}\big(\beta_{0},r_{\perp}\big)\subset\mathbb{R}^{n-q}$ as shorthand notation to denote open balls centered at the critical point. Finally, we define two scalar quantities as the maximum norm of $\xi_{\parallel}(\alpha,p),\xi_{\perp}(\alpha,\beta,p)$ over open balls centered around the critical point, and state the main result from [gupta2024estimates]
$L_{\parallel}(r_{\parallel})$ $:=\sup\big\{\left\lVert{\xi_{\parallel}(\alpha,p)}\right\rVert:(\alpha,p)\in\mathcal{B}_{\parallel}(r_{\parallel})\big\},$ (8a)
$L_{\perp}(r_{\parallel},r_{\perp})$ $:=\sup\big\{\left\lVert{\xi_{\perp}(\alpha,\beta,p)}\right\rVert:(\alpha,p)\in\mathcal{B}_{\parallel}(r_{\parallel})$
 $\hskip 85.35826pt\text{ and }\beta\in\mathcal{B}_{\perp}(r_{\perp})\big\}.$ (8b)
Proposition II.3
[gupta2024estimates, Theorem 3.2] For all $r_{\parallel},r_{\perp}>0$ simultaneously satisfying
@@ III Simplified General LS Bounds
Before stating our main results in the next section, we present some results simplifying the statement of the general bounds of II.3. First we streamline the expressions (6a), (6b) in terms of the projections $P:=P_{\operatorname{range}(J^{\star})}=WW^{\top}$ and $Q:=P_{\ker(J^{\star})^{\perp}}=\bar{V}\bar{V}^{\top}$, and the partial derivatives of $\Phi$ at $(x^{\star},p^{\star})$ w.r.t. $x$ and $p$.
@@
$M_{\parallel}=\left\lVert{P_{\perp}D_{p}\Phi(x^{\star},p^{\star})}\right\rVert$
@@
From (6a) and definition of the spectral norm,
[…]
$=\rho\Bigl(\bigl(D_{p}\Phi(x^{\star},p^{\star})\bigr)^{\top}WW^{\top}D_{p}\Phi(x^{\star},p^{\star})\Bigr)$
$=\left\lVert{W^{\top}D_{p}\Phi(x^{\star},p^{\star})}\right\rVert^{2}.$
Taking the square root on both sides and applying II.1 completes the proof. ∎
@@
From its definition in (6b),
$M_{\perp}$ $=\left\lVert{\bigl(W^{\top}D_{x}\Phi(x^{\star},p^{\star})\bar{V}\bigr)^{-1}}\right\rVert$
$=\sigma_{\max}\left(\bigl(W^{\top}J^{\star}\bar{V}\bigr)^{-1}\right)$
$=\Bigl[\sigma_{\min}\left(\Sigma\right)\Bigr]^{-1}=1/\sigma_{\min}(J^{\star}).$
since $\Sigma$ from (1) is the diagonal matrix of all non-zero singular values of $J^{\star}$. ∎
Next, we prove that the second inequality condition in the originally derived II.3 is redundant, and state a simplified version of the bounds.
@@
there exists a smooth implicit map $\varphi:\mathcal{B}_{\parallel}(r_{\parallel})\to\mathcal{B}_{\perp}(r_{\perp})$ that satisfies (5), where the quantities in (10) are defined by III.1, III.2, and equations (8a), (8b).
@@
$M_{\perp}L_{\perp}(r_{\parallel},r_{\perp})+M_{\perp}\big(M_{\parallel}+L_{\parallel}(r_{\parallel})\big)\frac{r_{\parallel}}{r_{\perp}}<1.$ (11)
Observe that all the quantities on the left hand side of (11) are nonnegative scalars. Therefore by monotonicity of addition, it must necessarily hold that $M_{\perp}L_{\perp}(r_{\parallel},r_{\perp})<1$, i.e. (9b) is satisfied. The theorem statement then follows directly from II.3. ∎
@@ IV Main Results: Network Bounds
where $x\in\mathbb{R}^{n}$ is the state vector, $p\in\mathbb{R}$ is a parameter, $b\in\mathbb{R}^{n}$ is a constant input or bias, $A\in\mathbb{R}^{n\times n}$, and $C=\operatorname{diag}(c_{1},\dots,c_{n})$ with $c_{i}\in\mathbb{R}^{>0}$. The function $S(z)=\begin{bmatrix}S_{1}(z_{1})&\dots&S_{n}(z_{n})\end{bmatrix}^{\top}$ applies an element-wise smooth activation function to each input, and its derivative is denoted as $D_{z}S(z)=\operatorname{diag}\bigl(S^{\prime}_{1}(z_{1}),\dots,S^{\prime}_{n}(z_{n})\bigr)$. We will refer to (12) as a Hopfield or voltage model as it follows the structure of classic and modern associative memory networks [hopfield1984neurons, betteti2025input]. Second we will consider systems of the form
@@
with all the quantities defined the same as for the Hopfield models. We will refer to (13) as a firing rate model [wilson1972excitatory, betteti2025firing]. It has been shown that (12), (13) are related through changes of coordinates and time-varying input transformations [miller2012mathematical].
@@ IV-A Hopfield models
and projection matrices $P:=P_{\operatorname{range}(J^{\star}_{H})}=W_{H}W_{H}^{\top}$ and $Q:=P_{\ker(J^{\star}_{H})^{\perp}}=\bar{V}_{H}\bar{V}_{H}^{\top}$ from $J^{\star}_{H}$ as described in (1), (2) and (3). Next, we make the following assumptions:
@@
    The pair $(x^{\star},p^{\star})$ is an equilibrium of (12), i.e.,
    $\Phi(x^{\star},p^{\star})=-Cx^{\star}+p^{\star}AS(x^{\star})+b=0.$ (15)
@@
    The Jacobian $J^{\star}_{H}$ has a zero eigenvalue with algebraic and geometric multiplicity $q$, and $\mathrm{Re}\,(\lambda_{j})\neq 0$ for all other eigenvalues.
To apply Theorem III.3 to (12), we must compute the various quantities in the inequality condition (10). In the following sequence of lemmas we derive expressions needed to define $L_{\parallel}(r_{\parallel}),L_{\perp}(r_{\parallel},r_{\perp})$ in (10), exploiting the form of Hopfield dynamics to express these quantities in terms of the network matrix $A$, activation function $S$, and its first derivatives.
@@
$\left\lVert{\xi_{\parallel}(\alpha,p)}\right\rVert=\left\lVert{PA\begin{bmatrix}\Delta_{\parallel}(\alpha,p)V&\bar{S}_{\parallel}(\alpha)\end{bmatrix}}\right\rVert$ where $\Delta_{\parallel}(\alpha,p)$ and $\bar{S}_{\parallel}(\alpha)$ are defined as
@@
$\bigl(-C+p^{\star}AD_{z}S(x^{\star})\bigr)V=0\implies CV=p^{\star}AD_{z}S(x^{\star})V,$
which allows us to simplify $\xi_{\parallel\alpha}(\alpha,p)$ as
@@
$=A\Delta_{\parallel}(\alpha,p)V$
Next, we simplify $\xi_{\parallel p}(\alpha,p)$ as
$\xi_{\parallel p}(\alpha,p)$ $:=D_{p}\Phi\bigl(\Gamma(\alpha,\beta_{0}),p\bigr)-D_{p}\Phi(x^{\star},p^{\star})$
$=A\Bigl(S\bigl(\Gamma(\alpha,\beta_{0})\bigr)-S(x^{\star})\Bigr)=A\bar{S}_{\parallel}(\alpha).$
Finally, combining our expressions for $\xi_{\parallel\alpha}$ and $\xi_{\parallel p}$ yields
$\xi_{\parallel}(\alpha,p)=W^{\top}A\begin{bmatrix}\Delta_{\parallel}(\alpha,p)V&\bar{S}_{\parallel}(\alpha)\end{bmatrix}$.
The desired expression for $\left\lVert{\xi_{\parallel}(\alpha,p)}\right\rVert$ follows directly by applying II.1 with to the above. ∎
@@
$\Delta_{\perp}(\alpha,\beta,p):=pD_{z}S\bigl(\Gamma(\alpha,\beta)\bigr)-p^{\star}D_{z}S(x^{\star}).$ (17)
@@
$\xi_{\perp}(\alpha,\beta,p)$ $=W^{\top}\Bigl(D_{x}\Phi\bigl(\Gamma(\alpha,\beta),p\bigr)-D_{x}(x^{\star},p^{\star})\Bigr)\bar{V}$
$=W^{\top}\Bigl(pAD_{z}S\bigl(\Gamma(\alpha,\beta)\bigr)-p^{\star}AD_{z}S(x^{\star})\Bigr)\bar{V}$
$=W^{\top}A\Delta_{\perp}(\alpha,\beta,p)\bar{V}.$
The desired expression follows directly by applying II.1 and II.2 in succession as follows: $\left\lVert{W^{\top}A\Delta_{\perp}(\alpha,\beta,p)\bar{V}}\right\rVert=\left\lVert{WW^{\top}A\Delta_{\perp}(\alpha,\beta,p)\bar{V}\bar{V}^{\top}}\right\rVert=\left\lVert{PA\Delta_{\perp}(\alpha,\beta,p)Q}\right\rVert$
With these lemmas established, we state the first result.
Theorem IV.3 (Bounds of bifurcation in Hopfield models)
Consider (12) under 1. For all $r_{\parallel},r_{\perp}>0$ satisfying
$L_{\parallel}(r_{\parallel})r_{\parallel}+L_{\perp}(r_{\parallel},r_{\perp})r_{\perp}\\ <\sigma_{\min}\left(J^{\star}_{H}\right)r_{\perp}-\tfrac{1}{p^{\star}}\left\lVert{P(Cx^{\star}-b)}\right\rVert r_{\parallel}$ (18)
@@
$L_{\parallel}(r_{\parallel})$ $=\sup\Bigl\{\left\lVert{PA\begin{bmatrix}\Delta_{\parallel}(\alpha,p)V&\bar{S}_{\parallel}(\alpha)\end{bmatrix}}\right\rVert$
$\hskip 51.21495pt:(\alpha,p)\in\mathcal{B}\bigl((\alpha_{0},p^{\star}),r_{\parallel}\bigr)\Bigr\},$
$L_{\perp}(r_{\parallel},r_{\perp})$ $=\sup\Bigl\{\left\lVert{PA\Delta_{\perp}(\alpha,\beta,p)Q}\right\rVert$
$:(\alpha,p)\in\mathcal{B}\bigl((\alpha_{0},p^{\star}),r_{\parallel}\bigr),\;\beta\in\mathcal{B}(\beta_{0},r_{\perp})\Bigr\},$
with $\Delta_{\parallel},\Delta_{\perp},\bar{S}_{\parallel}$ defined in (16), (17), there exists a smooth implicit map $\varphi:\mathcal{B}\bigl((\alpha_{0},p^{\star}),r_{\parallel}\bigr)\to\mathcal{B}(\beta_{0},r_{\perp})$ satisfying (5).
@@
From the equilibrium condition (15) we see that $AS(x^{\star})=\frac{1}{p^{\star}}(Cx^{\star}-b)$. The statement then immediately follows from applying III.1 to obtain
$M_{\parallel}=\left\lVert{PAS(x^{\star})}\right\rVert=\left\lVert{\tfrac{1}{p^{\star}}P(Cx^{\star}-b)}\right\rVert.$ (20)
Substituting the values of (6b), (8a), (8b) from III.2, IV.1, IV.2 into (10) from Theorem III.3 yields the theorem statement. ∎

### v3

@@ II-C Bounds of Validity for Lyapunov–Schmidt reduction
A key limitation of Lyapunov–Schmidt reduction is that the equivalence between the reduced bifurcation equations and the equilibria of the original system holds only locally, restricting the validity of analysis based on this reduction to a neighbourhood of the singular point. Quantifying this neighbourhood is crucial for evaluating the robustness of conclusions drawn from the reduced equations. To address this limitation, in previous work we derived a general bound of validity for Lyapunov–Schmidt reduction for bifurcations of equilibria in systems of ordinary differential equations [13]. In this section we summarize our recent result.
In the context of Lyapunov–Schmidt reduction, we derive explicit lower bounds on the neighbourhoods of the implicit map $\varphi$ as open balls, i.e. $\mathcal{B}\big((\alpha_{0},p^{\star}),r_{\parallel}\big)\subset\mathcal{N}\big((\alpha_{0},p^{\star})\big)$ and $\mathcal{B}(\beta_{0},r_{\perp})\subset\mathcal{N}(\beta_{0})$. The radii $r_{\parallel}$ and $r_{\perp}$ quantify the neighbourhood of the singular point $(\alpha,\beta,p)=(\alpha_{0},\beta_{0},p^{\star})$ within which the reduced and full bifurcation diagrams coincide topologically: $r_{\parallel}$ specifies the radius along the critical subspace $\ker(J^{\star})\times\Lambda$, while $r_{\perp}$ specifies the corresponding radius along $\ker(J^{\star})^{\perp}$. The algorithm to explicitly determine the bounding radii $r_{\parallel}$ and $r_{\perp}$ is described below.
@@
$M_{\perp}$ $:=\left\lVert{\left(W^{\top}J^{\star}\overline{V}\right)^{-1}}\right\rVert.$ (5b)
@@
$\xi_{\perp}(\alpha,\beta,p)$ $:=W^{\top}\bigl(D_{x}\Phi\bigl(\Gamma(\alpha,\beta),p\bigr)-J^{\star}\bigr)W,$ (6b)
where $\xi_{\parallel p}(\alpha,p)=D_{p}\Phi\bigl(\Gamma(\alpha,\beta_{0}),p\bigr)-D_{p}\Phi(x^{\star},p^{\star})$ and $\xi_{\parallel\alpha}(\alpha,p)=D_{x}\Phi\bigl(\Gamma(\alpha,\beta_{0}),p\bigr)V$. The maps (6) quantify first-order variations of the Jacobian $D_{x}\Phi$ and parameter derivative $D_{p}\Phi$ along the critical subspace and its orthogonal complement. We use $\mathcal{B}_{\parallel}(r_{\parallel}):=\mathcal{B}\big((\alpha_{0},p^{\star}),r_{\parallel}\big)\subset\mathbb{R}^{q}\times\Lambda$ and $\mathcal{B}_{\perp}(r_{\perp}):=\mathcal{B}\big(\beta_{0},r_{\perp}\big)\subset\mathbb{R}^{n-q}$ as shorthand notation to denote open balls centred at the critical point. Finally, we define two scalar quantities as the maximum norm of $\xi_{\parallel}(\alpha,p),\xi_{\perp}(\alpha,\beta,p)$ over open balls centred around the critical point, and state the main result from [13]
$L_{\parallel}(r_{\parallel})$ $:=\sup_{(\alpha,p)\in\mathcal{B}_{\parallel}(r_{\parallel})}\left\lVert{\xi_{\parallel}(\alpha,p)}\right\rVert$ (7a)
$L_{\perp}(r_{\parallel},r_{\perp})$ $:=\sup_{\begin{subarray}{c}(\alpha,p)\in\mathcal{B}_{\parallel}(r_{\parallel})\\ \beta\in\mathcal{B}_{\perp}(r_{\perp})\end{subarray}}\left\lVert{\xi_{\perp}(\alpha,\beta,p)}\right\rVert.$ (7b)
Theorem II.2
[13, Theorem 3.2] For all $r_{\parallel},r_{\perp}>0$ simultaneously satisfying
@@ III Simplified General LS Bounds
Before stating our main results in the next section, we present some results simplifying the statement of the general bounds of Theorem II.2. First we streamline the expressions (5a), (5b) in terms of the projections $P_{\operatorname{range}(J^{\star})}=WW^{\top}$ and the partial derivatives of $\Phi$ at $(x^{\star},p^{\star})$ w.r.t. $x$ and $p$.
@@
$M_{\parallel}=\left\lVert{P_{\operatorname{range}(J^{\star})}D_{p}\Phi(x^{\star},p^{\star})}\right\rVert.$
@@
From (5a) and definition of spectral radius, $M_{\parallel}^{2}=\\ \rho\left(\begin{bmatrix}\mathbf{0}_{n-q}\!&\!W^{\top}J_{p}^{\star}\end{bmatrix}^{\top}\begin{bmatrix}\mathbf{0}_{n-q}\!&\!W^{\top}J_{p}^{\star}\end{bmatrix}\right)=\rho\bigl(J_{p}^{\star\top}WW^{\top}J_{p}^{\star}\bigr)\\ =\left\lVert{W^{\top}J_{p}^{\star}}\right\rVert^{2},$ where $J_{p}^{\star}:=D_{p}\Phi(x^{\star},p^{\star})$. Taking square root and applying Lemma II.1 completes the proof. ∎
@@
From its definition in (5b), $M_{\perp}=\left\lVert{\bigl(W^{\top}J^{\star}\overline{V}\bigr)^{-1}}\right\rVert=\left\lVert{\bigl(W^{\top}W\Sigma\overline{V}^{\top}\overline{V}\bigr)^{-1}}\right\rVert=\left\lVert{\Sigma^{-1}}\right\rVert=\sigma_{\max}(\Sigma^{-1})=\frac{1}{\sigma_{\min}(\Sigma)}\\ =1/\sigma_{\min}(J^{\star}),$ since $\Sigma$ is the diagonal matrix of singular values (all strictly positive) of $J^{\star}$. ∎
Next, we prove that the second inequality condition in the originally derived Theorem II.2 is redundant, and state a simplified version of the bounds.
@@
there exists a smooth implicit map $\varphi:\mathcal{B}_{\parallel}(r_{\parallel})\to\mathcal{B}_{\perp}(r_{\perp})$ that satisfies (4), where the quantities in (9) are defined by Lemma III.1, Lemma III.2, and equations (7a), (7b).
@@
$M_{\perp}L_{\perp}(r_{\parallel},r_{\perp})+M_{\perp}\big(M_{\parallel}+L_{\parallel}(r_{\parallel})\big)\frac{r_{\parallel}}{r_{\perp}}<1.$
Since all the quantities on the above left hand side are non-negative scalars, it follows that $M_{\perp}L_{\perp}(r_{\parallel},r_{\perp})<1$ must necessarily hold, i.e. (8b) is satisfied. The theorem statement then follows directly from Theorem II.2. ∎
@@ IV Main Results: Network Bounds
where $x\in\mathbb{R}^{n}$ is the state vector, $p\in\mathbb{R}$ is a parameter, $b\in\mathbb{R}^{n}$ is a constant input or bias, $A\in\mathbb{R}^{n\times n}$, and $C=\operatorname{diag}(c_{1},\dots,c_{n})$ with $c_{i}\in\mathbb{R}^{>0}$. The function $S(z)=\begin{bmatrix}S_{1}(z_{1})&\dots&S_{n}(z_{n})\end{bmatrix}^{\top}$ applies an element-wise smooth activation function to each input, and its derivative is denoted as $D_{z}S(z)=\operatorname{diag}\bigl(S^{\prime}_{1}(z_{1}),\dots,S^{\prime}_{n}(z_{n})\bigr)$. We will refer to (10) as a Hopfield or voltage model as it follows the structure of classic and modern associative memory networks [14, 3]. The second class we will consider are systems of the form
@@
with all the quantities defined the same as for the Hopfield models. We will refer to (11) as a firing rate model [23, 2]. It has been shown that (10), (11) are related through changes of coordinates and time-varying input transformations [19].
@@ IV-A Hopfield models
and projection matrices $P:=P_{\operatorname{range}(J^{\star}_{H})}=W_{H}W_{H}^{\top}$ and $Q:=P_{\ker(J^{\star}_{H})^{\perp}}=\overline{V}_{H}\overline{V}_{H}^{\top}$ from $J^{\star}_{H}$ as described in (1) and (2). Next, we make the following assumptions:
@@
    The pair $(x^{\star},p^{\star})$ is an equilibrium of (10).
@@
    The Jacobian $J^{\star}_{H}$ has a zero eigenvalue with multiplicity $q$, and $\mathrm{Re}\,(\lambda_{j})\neq 0$ for all other eigenvalues.
To apply Theorem III.3 to (10), we compute the various quantities in the inequality condition (9). In the following sequence of lemmas we derive expressions needed to express $L_{\parallel}(r_{\parallel}),L_{\perp}(r_{\parallel},r_{\perp})$ in (9), exploiting the form of Hopfield dynamics to express these quantities in terms of the network matrix $A$, activation function $S$, and its first derivatives.
@@
$\left\lVert{\xi_{\parallel}(\alpha,p)}\right\rVert=\left\lVert{PA\begin{bmatrix}\Delta_{\parallel}(\alpha,p)V&\bar{S}_{\parallel}(\alpha)\end{bmatrix}}\right\rVert$,
where $\Delta_{\parallel}(\alpha,p)$ and $\bar{S}_{\parallel}(\alpha)$ are defined as
@@
$\bigl(-C+p^{\star}AD_{z}S(x^{\star})\bigr)V=0\implies CV=p^{\star}AD_{z}S(x^{\star})V,$ which allows us to simplify $\xi_{\parallel\alpha}(\alpha,p)$ as
@@
$=:A\Delta_{\parallel}(\alpha,p)V$
Next, we express $\xi_{\parallel p}(\alpha,p)=D_{p}\Phi\bigl(\Gamma(\alpha,\beta_{0}),p\bigr)-\\ D_{p}\Phi(x^{\star},p^{\star})=A\bigl(S\bigl(\Gamma(\alpha,\beta_{0})\bigr)-S(x^{\star})\bigr)=A\bar{S}_{\parallel}(\alpha).$
Finally, combining our expressions for $\xi_{\parallel\alpha}$ and $\xi_{\parallel p}$ yields $\xi_{\parallel}(\alpha,p)=W^{\top}A\begin{bmatrix}\Delta_{\parallel}(\alpha,p)V&\bar{S}_{\parallel}(\alpha)\end{bmatrix}$. Applying Lemma II.1 to this yields the desired expression. ∎
@@
$\Delta_{\perp}(\alpha,\beta,p):=pD_{z}S\bigl(\Gamma(\alpha,\beta)\bigr)-p^{\star}D_{z}S(x^{\star}).\vskip 5.69054pt$ (14)
@@
$\xi_{\perp}(\alpha,\beta,p)$ $=W^{\top}\Bigl(D_{x}\Phi\bigl(\Gamma(\alpha,\beta),p\bigr)-D_{x}\Phi(x^{\star},p^{\star})\Bigr)\overline{V}$
$=W^{\top}\Bigl(pAD_{z}S\bigl(\Gamma(\alpha,\beta)\bigr)-p^{\star}AD_{z}S(x^{\star})\Bigr)\overline{V}$
$=:W^{\top}A\Delta_{\perp}(\alpha,\beta,p)\overline{V}.$
The desired expression follows directly by applying Lemma II.1 as follows: $\left\lVert{W^{\top}A\Delta_{\perp}(\alpha,\beta,p)\overline{V}}\right\rVert=\\ \left\lVert{WW^{\top}A\Delta_{\perp}(\alpha,\beta,p)\overline{V}\,\overline{V}^{\top}}\right\rVert=\left\lVert{PA\Delta_{\perp}(\alpha,\beta,p)Q}\right\rVert.$
With these lemmas in place, we now state the bounds from Theorem III.3 for bifurcations in the Hopfield model.
Theorem IV.3
Consider the Hopfield dynamics (10) under 1. For all $r_{\parallel},r_{\perp}>0$ satisfying
$L_{\parallel}(r_{\parallel})r_{\parallel}+L_{\perp}(r_{\parallel},r_{\perp})r_{\perp}\\ <\sigma_{\min}\left(J^{\star}_{H}\right)r_{\perp}-\tfrac{1}{\left\lvert{p^{\star}}\right\rvert}\left\lVert{P(Cx^{\star}-b)}\right\rVert r_{\parallel}$ (15)
@@
$L_{\parallel}(r_{\parallel})$ $=\underset{(\alpha,p)\in\mathcal{B}_{\parallel}(r_{\parallel})}{\sup}\left\lVert{PA\begin{bmatrix}\Delta_{\parallel}(\alpha,p)V&\bar{S}_{\parallel}(\alpha)\end{bmatrix}}\right\rVert$
$L_{\perp}(r_{\parallel},r_{\perp})$ $=\underset{\begin{subarray}{c}(\alpha,p)\in\mathcal{B}_{\parallel}(r_{\parallel})\\ \beta\in\mathcal{B}_{\perp}(r_{\perp})\end{subarray}}{\sup}\left\lVert{PA\Delta_{\perp}(\alpha,\beta,p)Q}\right\rVert,$
with $\Delta_{\parallel}$, $\Delta_{\perp}$, $\bar{S}_{\parallel}$ defined in (13a), (14), there exists a smooth implicit map $\varphi:\mathcal{B}_{\parallel}(r_{\parallel})\to\mathcal{B}_{\perp}(r_{\perp})$ satisfying (4).
@@
From the equilibrium condition, we see that $AS(x^{\star})=\frac{1}{p^{\star}}(Cx^{\star}-b)$. The statement then immediately follows from applying Lemma III.1 to obtain
$M_{\parallel}=\left\lVert{PAS(x^{\star})}\right\rVert=\tfrac{1}{\left\lvert{p^{\star}}\right\rvert}\left\lVert{P(Cx^{\star}-b)}\right\rVert.$ (16)
Substituting (5b), (7a), (7b) from Lemmas III.2, IV.1, IV.2 into (9) from Theorem III.3 yields the theorem statement. ∎